OpenStudy (anonymous):

The equation to represent the cost of raising a child 18 years is C = 2.42N + 35,000. C is the cost and N represents the income of the parents. The Boyds have an income of $28,700. What is their cost of raising a child to 18 years of age? Answer $68,880 $69,454 $103,880 $104,454

OpenStudy (anonymous):

The equation to represent the cost of raising a child 18 years is C = 2.42N + 35,000. C is the cost and N represents the income of the parents. If the cost of raising a child to 18 years is $150,000, what is the income of the parents? (Hint: this time you will substitute for C and solve for N.) Round to the nearest dollar. Answer $47,521 $6,198 $35,295 $42,843

OpenStudy (anonymous):

Given C = 2.42N + 35,000 Substituting N = $28,700 in the above eq we find: \[\huge C = 2.42 \times 28700 + 35,000 \] \[\huge \rightarrow C=69454 +35000 =104454\] Hence the cost of raising a child to 18 years of age is 104454 dollars. @jashi34

OpenStudy (anonymous):

Thanks that helps alot

OpenStudy (anonymous):

Given: \[\huge C = 2.42N + 35,000.\] Substituting C = $150,000 in the above eq we find. \[\huge 150,000 = 2.42N + 35,000\] \[\huge 150,000 -35000= 2.42N \] \[\huge 115,000 = 2.42N \rightarrow N = \frac{115000}{2.42}\] \[N = 47520.66 \approx 47521\] Hence the income of the parent is 47521 dollars. @jashi34